OpenStudy (anonymous):

Find cos theta in (pi/2,pi) so that cot theta = -12/5

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

It is impossible for the value of cosine of any angle to be larger than 1 and smaller than -1. We notiuce that cos theta = -12/5 = -2.4 < -1, therefore there is no value for theta in (pi/2 , pi)

OpenStudy (anonymous):

Cot theta...

OpenStudy (anonymous):

\[\frac{\text{Sin}[\theta ]}{\text{Cos}[\theta ]}=\frac{-5}{12}\]

OpenStudy (anonymous):

Ok, that's tan theta, so what's cos theta?

OpenStudy (anonymous):

\[\frac{12}{-5} \text{Sin}[\theta ]=\text{Cos}[\theta ]\]

OpenStudy (anonymous):

I want a value...

OpenStudy (anonymous):

To find this value, you'll have to apply Pythagorean identity: 1 + (tan theta)^2 = 1/(cos theta)^2 (cos theta) = - sqrt [1/1 + (tan theta)^2]

OpenStudy (anonymous):

@giorgia looks promising, so what is it then?

OpenStudy (anonymous):

You'll plug the value for tan theta and you'll raise to square: (cos theta) = - sqrt [1/1 +144/25] cos theta = -sqrt169/25 cos theta = -13/5

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

oh, yes!

OpenStudy (anonymous):

12/13 ?

OpenStudy (anonymous):

Think the problem is us have sqrt(1/1...) instead of 1/sqrt(1+

OpenStudy (anonymous):

@imranmeah Close....check interval.

OpenStudy (anonymous):

Sign of cos in second quadrant?

OpenStudy (anonymous):

-12/13